OpenStudy (anonymous):

Part a.) Write the equation of the line that has a slope of -2 and a y-intercept of -6. Part b.) Write the equation of the line that passes through the point (6, -2) and has a slope of -2/3.

Directrix (directrix):

Y = mx + b Y = -2x + (-6)

OpenStudy (anonymous):

a . y = mx + b y = -2x -6 b. y - y1 = m(x - x1) y +2 = -2/3 (x -6)

OpenStudy (anonymous):

finish the last line of b by getting into slope intercept form

Directrix (directrix):

which yields y = (-2/3) x + 2